• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M. Cummings

    Log entry time 17:49:12 on March 09, 2012

    Entry number 365115

    This entry is a followup to: 365063

    keyword=T2 rate problem

    As seen in the previous log, this morning Ryan and I thought we had fixed
    the rate problem with T2.  However, after taking a cosmic run I found
    that there was a significant number of events (~1%) that corresponded to
    both T1 and T2; that is, DR.evtypebits==6.
    I took a closer look at the strobe signal compared to the s1 "or" signal,
    and found that the strobe was coming too late.  I also found that the
    strobe was more narrow on the RHRS (~40ns) compared to the LHRS (~80ns).
    This afternoon Vince and I took a detailed look at the timing of the LHRS
    and found that the strobe signal came ~64ns after the logic signals of
    s1, s2m and the Cherenkov.  On the RHRS we first moved the strobe signal
    from a 754 logic unit to a 755 logic unit, which allowed for the signal
    to be made ~80ns wide.  We also found that the delay which had been added
    to the strobe signal this morning was much more than was necessary.  For
    some reason the programmable delay module in slot 21 of crate 1 adds much
    more delay than the module is actually programmed for.  We moved the
    strobe to a different delay module and added an appropriate amount of
    delay so that the strobe came between 30-80ns after the logic signals of
    s1, s2m and the Cherenkov.
    The T2 rate is now ~1-2 Hz for cosmics, comparable to what we found in
    December. There is also no longer an overlap between T1 and T2.  We will
    continue to monitor when the beam comes back.

    A copy of this log entry has been emailed to: vasulk, rbziel, melissac